OpenStudy (anonymous):

Medal to most helpful. Simplify 2 times the 4th root of 80. Thanks.

OpenStudy (anonymous):

\[2 \sqrt[4]{80}\]

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle 2\cdot \sqrt[4]{80} }\)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle 2\cdot \sqrt[4]{80} }\) \(\large\color{#000000 }{ \displaystyle 2\cdot \sqrt[4]{16\times5} }\) \(\large\color{#000000 }{ \displaystyle 2\cdot \sqrt[4]{16}\times\sqrt[4]{5} }\)

OpenStudy (solomonzelman):

Can you go from there?

OpenStudy (anonymous):

Could you continue from there, please? @SolomonZelman

OpenStudy (solomonzelman):

Do you know what "forth root" means?

OpenStudy (anonymous):

one fourth power

OpenStudy (anonymous):

2 * 16^1/4 * 5^1/4 ??

OpenStudy (solomonzelman):

But that second component is integer when simplified.

OpenStudy (solomonzelman):

I will try to make the definition clear through one example. Just stay with me please

OpenStudy (solomonzelman):

You know that, \(\large\color{#000000 }{ \displaystyle 3\times 3\times 3\times 3 =81 }\) Ok?

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

Ok, good, and how many times did I multiply the 3?

OpenStudy (anonymous):

4 times

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \underbrace{3\times 3\times 3\times 3 }_{4~\rm times}=81 }\)

OpenStudy (anonymous):

I'm following.

OpenStudy (solomonzelman):

And that would mean \(\large\color{#000000 }{ \displaystyle 3^4=81 }\)

OpenStudy (solomonzelman):

And the inverse operation of \(\large\color{#000000 }{ \displaystyle 3^4=81 }\), is, \(\Large\color{#000000 }{ \displaystyle \sqrt[\color{red}{4}]{81} =3}\)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \sqrt[4]{81}=3 }\) and/because \(\large\color{#000000 }{ \displaystyle 3^4=81 }\) So, this way, \(\large\color{#000000 }{ \displaystyle \sqrt[4]{16}={\bf \quad is~?} }\)

OpenStudy (solomonzelman):

(try dividing 16 by 2 a bunch of times and tell me that 2^what? = 16)

OpenStudy (anonymous):

Answer Choices: (Just for reference) A. \[4\sqrt[4]{5}\] B. \[8\sqrt[4]{5}\] C. \[16\sqrt[4]{5}\] D. \[32\sqrt[4]{5}\]

OpenStudy (solomonzelman):

(This way you would be able to determine the 4th root of 16)

OpenStudy (anonymous):

2?

OpenStudy (solomonzelman):

Explain your question

OpenStudy (anonymous):

sorry it took so long i had to refresh openstudy

OpenStudy (anonymous):

the answer to your question is 2. the fourth root of 16

OpenStudy (solomonzelman):

Yes, the 4th root of 16 is 2.

OpenStudy (solomonzelman):

So let's come back to our problem.

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle 2\cdot \sqrt[4]{80} }\) \(\large\color{#000000 }{ \displaystyle 2\cdot \sqrt[4]{16\times5} }\) \(\large\color{#000000 }{ \displaystyle 2\cdot \sqrt[4]{16}\times\sqrt[4]{5} }\) this is where we left.

OpenStudy (anonymous):

yep

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle 2\cdot \color{red}{2}\times\sqrt[4]{5} }\)

OpenStudy (solomonzelman):

So your most simplified answer is?

OpenStudy (solomonzelman):

(Note that there is NO way to simplify \(\large\color{#000000 }{ \sqrt[4]{5} }\))

OpenStudy (anonymous):

\[4\sqrt[4]{5}\]

OpenStudy (solomonzelman):

Yup, that's exactly right!

OpenStudy (anonymous):

Okay, thank you for your help! I appreciate it. Cheers!

OpenStudy (solomonzelman):

YW